OpenStudy (anonymous):

GIVING MEDAL find the area of the triangle with a=4ft, b=8ft and c=11ft a. 12.3ft^2 b. 9.3ft^2 c. 14.3ft^2 d. 13.3ft^2

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=-YI6UC4qVEY

OpenStudy (campbell_st):

you need heron's formula for the area of a triangle when given 3 sides you need to find a value s which is \[s = \frac{a + b+c}{2}\] you have a = 4, b = 8 and c = 11 then after finding s the area is \[A = \sqrt{s \times(s -a)\times(s -b) \times (s -c)}\] just substitute and the calculate the answer.

OpenStudy (campbell_st):

here is a link to explain it http://www.mathopenref.com/heronsformula.html

OpenStudy (anonymous):

i never learned that